the sum of 3 consecutive even integers is fourteen less than twice the smallest of these intergers
Find the smallest and greatest interger

Answer by geetha_rama(94)

Let X be the smallest integer
X + X+2 +X+4 = 2X - 14 (Equation 1) (Sum of the three integers is 14 less than twice the smallest of these integers)
3X + 6 = 2X -14
3X - 2X = -14-6
X = -20
Answer : Three integers are -20, -18, -16
LHS of Eq 1 : Sum of these three integrers = -54
RHS : 2X - 14 = 2*-20 -14 = -54